Hannah and Freancine have $120.00. Hannah and Peter have $230.00. Peter has 6 times as much money as Freancine. How much money does Hannah have? (real-world problems: Multiplication and Division)
Answer by robertb(5830)

h +f = 120
h + p = 230
p = 6f
Substitute the 3rd equation into the 2nd equation: h + 6f = 230. Then from this, h = 230-6f. Put this into the 1st equation: 230-6f + f = 120, or
-5f = -110, or f = 22. This means h = 120-f = 120-22 = 98. Hannah has $98.